OpenStudy (anonymous):

what is -3/8 divided by 7/12

OpenStudy (haleyelizabeth2017):

\[-\frac{3}{8}*\frac{12}{7}\] is the same thing.

OpenStudy (haleyelizabeth2017):

multiply 3 by 12, and 8 by 7. what is 3*12?

OpenStudy (mathmath333):

=-9/14

OpenStudy (kweenlexis):

yeah what @mathmath333 said